OpenStudy (anonymous):

Which represents the inverse of the function f(x) = 4x? h(x) = x + 4 h(x) = x – 4 h(x) = x3/4x h(x) = x1/4x

OpenStudy (campbell_st):

to find an inverse, swap x and y x = 4y now make y the subject.... and the answer is the inverse

OpenStudy (anonymous):

which is 3????

OpenStudy (anonymous):

\[given\ f(x)= 4x\\to \ find \ inverse\ of\ f(x) \ say \ h(x)\\ switch\ x \ and \ f(x) \\ i.e.\ x=4f(x)\\solve\ for\ f(x)\\here\ f(x)=\frac{ x }{ 4 }\\replace \ f(x)\ by\ h(x)\\ get\ h(x)= \frac{ x }{ 4 }\]

OpenStudy (anonymous):

@elzabeth22950 sure! options 4th posted correctly? \[h(x)=\frac{ x }{ 4} is \ correct.\]